# Exotic synchronization in continuous time crystals outside the symmetric subspace

### v1

@@ Appendix C Appendix C: Cluster synchronization
We will now discuss the effect of small changes in the initial state on the synchronization of CTCs. As an example, we consider a change in the initial states for CTCs belonging to ensemble $\varepsilon_{2}$ such that the distribution of $\{|\lambda_{1}^{\alpha}|\}$ is from $0.75$ to $0.8$. The choice of initial states for the uncoupled CTCs is shown in Figs. 5(a,d). Now we study the effect of choosing this distribution on the synchronization of CTCs and we only consider all-to-all coupling in this case. This small change results in the emergence of a new synchronization phase, termed cluster synchronization as shown in Figs. 5(b,e). As the name suggests, there exist two clusters of synchronized CTCs with different frequencies. Further increasing the coupling strength to $\Gamma/\kappa=1.25$ results in complete synchronization and these two clusters oscillate at the same frequency as shown in Figs. 5(c,f). These findings demonstrate that different choices of initial states can result in different regimes of synchronization.
@@
Figure 5: Here we consider $n=20$ CTCs having the same $\Omega_{\alpha}/\kappa=0.9$ which are initialized with different $m$ subspaces. Subfigure (a,d) represents uncoupled time crystals. For finite coupling strength $\Gamma/\kappa=0.35$, there are two clusters of synchronized CTCs with different frequencies as shown in (b,e). At a coupling strength $\Gamma/\kappa=1.25$, all the CTCs get synchronized as shown in (c,f).

### v2

@@ Appendix D Cluster synchronization
We will now discuss the effect of small changes in the initial state on the synchronization of CTCs. As an example, we consider a change in the initial states for CTCs belonging to ensemble $\varepsilon_{2}$ such that $\{|\Lambda_{1}^{\alpha}|\}$ are uniformly sampled in the interval from $0.75$ to $0.8$. The choice of initial states for the uncoupled CTCs is shown in Figs. S3(a,d). Now we study the effect of choosing this distribution on the synchronization of CTCs and we only consider all-to-all coupling in this case. This small change results in the emergence of a new synchronization phase, termed cluster synchronization as shown in Figs. S3(b,e). As the name suggests, there exist two clusters of synchronized CTCs with different frequencies. Further increasing the coupling strength to $\Gamma/\kappa=1.25$ results in complete synchronization and these two clusters oscillate at the same frequency as shown in Figs. S3(c,f). These findings demonstrate that different choices of initial states can result in different regimes of synchronization.
@@
Figure S3: Here we consider $n=20$ CTCs having the same $\Omega_{\alpha}/\kappa=0.9$ which are initialized with different $m$ subspaces. Subfigure (a,d) represents uncoupled CTCs. The frequencies of the uncoupled CTCs are uniformly sampled in the window from $0.2$ to $0.25$ in ensemble $\varepsilon_{1}$ and from $0.75$ to $0.8$ in ensemble $\varepsilon_{2}$. For finite coupling strength $\Gamma/\kappa=0.35$, there are two clusters of synchronized CTCs with different frequencies, as shown in (b,e). At a coupling strength $\Gamma/\kappa=1.25$, all the CTCs get synchronized as shown in (c,f).
Appendix E Characterization of different phases using Pearson coefficient and maximum Lyapunov exponent
Here, we provide a detailed characterization of the various synchronization regimes discussed in the main text. We consider an ensemble of $n=100$ CTCs that are divided into two subgroups $\varepsilon_{1,2}$ containing $50$ CTCs each. The frequencies of the CTCs in each subgroup $\varepsilon_{1,2}$ are sampled from Gaussian distributions with distinct mean frequencies $\bar{\omega}_{1,2}$ and standard deviations $\sigma_{1,2}$. As in Fig. 4(e,f) of the main text, we use $\sigma_{1}=\sigma_{2}=0.1$, and vary $\Delta=\bar{\omega}_{1}-\bar{\omega}_{2}$ and $\Gamma$ to explore six regimes of synchronization marked by different symbols in Figs. S4 and S5. We study the average value of the Pearson coefficient $\bar{\mathcal{C}}$ and the maximum Lyapunov exponent $\lambda_{L}$ defined in the main text to characterize these regimes.
[…]
(pink)’($\Gamma=0.8,\Delta=0.3$), ‘(yellow)’($\Gamma=0.8,\Delta=0.2$) and $\blacktriangle$(black)’($\Gamma=0.8,\Delta=0.1$) in subfigure (a).
Partial Synchronization —
[…]
  • Chimera state: For $\Gamma=0.35$ and $\Delta=0.6$ (denoted by ‘’), there is only one prominent cluster of synchronized CTCs while the other CTCs remain unsynchronized. This can be understood from Fig. S4(c), where the CTCs of subgroup $\varepsilon_{1}$ show a high amount of correlation while the remainder of the CTCs are almost uncorrelated. Therefore, the CTCs of subgroup $\varepsilon_{1}$ have a common frequency, while the other CTCs oscillate with different frequencies as shown in Fig. S5(c). Thus, the system exhibits a chimera state.
  • Chimera state with partial oscillation death: We next discuss a case in which the Pearson correlation matrix looks qualitatively similar to that of a chimera state (see Fig. S4(c)). This can be observed for $\Gamma=0.8,\Delta=0.6$ (denoted by $\bullet$’), as shown in Fig. S4(d). Note that $\lambda_{L}<0$, see Fig. S5(a). Thus, some of the CTCs cease to oscillate leading to partial oscillation death. The same conclusion follows from Fig. S5(d) where all the CTCs of subgroup $\varepsilon_{1}$ have zero frequency and hence melt down. The steady state of the melted CTCs corresponds to highly correlated fixed points, and the correlation matrix exhibits a block diagonal structure similar to what was observed for chimera states.
As we have seen, $\bar{\mathcal{C}}$ alone does not provide sufficient information to distinguish different cases of partial synchronizations. Therefore, we investigated the correlation matrix and the maximum Lyapunov exponent to differentiate between these cases.
Oscillation death —
We now discuss the case of complete oscillation death, where all the CTCs melt down and stop oscillating. This can be observed for $\Gamma=0.8$ and $\Delta=0.3$ (represented by ‘ ’), where $\lambda_{L}\ll 0$, as shown in Fig. S5(a). Indeed, from Fig. S5(e), we can observe that all the CTCs have zero frequency. Therefore the system settles down to a fixed point corresponding to complete oscillation death. The Pearson coefficient matrix shows a large amount of correlation between all the CTCs.
Mean-field chaos —
By further reducing the detuning to $\Delta=0.2$ while keeping $\Gamma=0.8$ (represented by ‘’), the mean-field dynamics becomes chaotic, indicated by a positive value of $\lambda_{L}$ in Fig. S5(a). This leads to a lower value of $\bar{\mathcal{C}}$ as shown in Fig. S4(a). Although the correlation matrix exhibits a reduced amount of correlations between the CTCs due to the chaotic nature of the mean-field dynamics, the finite value of its off-diagonal elements indicates the presence of chaotic synchronization.
Complete synchronization —
Finally, for small detuning $\Delta=0.1$ and $\Gamma=0.8$ (represented by $\blacktriangle$’), the system becomes completely synchronized such that $\bar{\mathcal{C}}\approx 1$ and $\lambda_{L}\approx 0$. This regime is indicated by Fig. S4(g), where all the off-diagonal elements of $\mathcal{C}$ are non-zero, representing a large amount of correlation between all CTCs. The complete synchronization can also be observed from Fig. S5(g), where all CTCs oscillate with a common frequency.
[…]
(pink)’($\Gamma=0.8,\Delta=0.3$), ‘(yellow)’($\Gamma=0.8,\Delta=0.2$) and $\blacktriangle$(black)’($\Gamma=0.8,\Delta=0.1$) in subfigure (a). All the other parameters are as in Fig. S4.